OpenStudy (anonymous):

Find (d/dx)(x^2)(e^(ln*x^3)). The answer is 5x^4.

OpenStudy (anonymous):

e^(ln x^3) = x^3 so x^2 * x^3 = x^5 and the integral of that is 5x^4

OpenStudy (anonymous):

HAHAHA, it was that easy? Thank you so much!

OpenStudy (anonymous):

i meant *derivative*

OpenStudy (anonymous):

Thanks a lot.